the perimeter of a rectangle is 41 inches and its area is 91 square inches. find the length of its shortest side.
Answer by amoresroy(361)

the perimeter of a rectangle is 41 inches and its area is 91 square inches. find the length of its shortest side.
Let x & y represent the dimensions of the rectangle
2x + 2y = 41
xy = 91
_________________
Substitute y = 91/x in 1st equation
2x + 2(91/x) = 41
Multiply both sides by x
2x^2+182+-41x = 0
Solve for x by factoring
(2x-13)(x-14)
2x-13=0
x = 6.5
The shortest side of the rectangle is 6.5 inches